# Conditioning LLMs to Generate Code-Switched Text: A Methodology Grounded in Naturally Occurring Data

### v1

@@ Abstract
Code-switching (CS) is still a critical challenge in Natural Language Processing (NLP). Current Large Language Models (LLMs) struggle to interpret and generate code-switched text, primarily due to the scarcity of large-scale CS datasets for training. This paper presents a novel methodology to generate CS data using LLMs, and test it on the English-Spanish language pair. We propose back-translating natural CS sentences into monolingual English, and using the resulting parallel corpus to fine-tune LLMs to turn monolingual sentences into CS. Unlike previous approaches to CS generation, our methodology uses natural CS data as a starting point, allowing models to learn its natural distribution beyond grammatical patterns. We thoroughly analyse the models’ performance through a study on human preferences, a qualitative error analysis and an evaluation with popular automatic metrics. Results show that our methodology generates fluent code-switched text, expanding research opportunities in CS communication, and that traditional metrics do not correlate with human judgement when assessing the quality of the generated CS data. We release our code and generated dataset under a CC-BY-NC-SA license. [^1] GitHub [Uncaptioned image]      HuggingFace[Uncaptioned image]
@@ 1 Introduction
Despite the prevalence of code-switching, most research in Natural Language Processing (NLP) assumes monolingualism as a standard for human communication. However, this implicit decision means that state-of-the-art models are not able to properly interpret or generate CS data. Even advances in multilingual language modelling (Lin et al., 2022; Chowdhery et al., 2023) have not led to significant improvements, and performance on CS data is still poor compared to performance on monolingual data (Aguilar et al., 2020; Winata et al., 2021). This occurs because there is little CS text available in the multilingual pretraining data. Similarly, there are no parallel datasets available to learn to generate CS in a supervised fashion, as one would expect for tasks such as machine translation. Finally, there is no clear methodology for evaluating automatically generated CS text, as it has specific needs different from other text generation tasks.
It is therefore crucial to develop methodologies to enable models to generate natural CS text and simultaneously implement robust evaluation frameworks that can assess how well NLP systems handle CS across multiple tasks. We argue that both of these goals require models that can conditionally generate CS from monolingual text. Consequently, our research focuses on the development of a methodology to fine-tune and evaluate LLMs on the task of CS generation, following two main research questions:
@@
Is it viable to approach CS generation with natural CS text as a starting point?
@@
Do automatic metrics for Natural Language Generation (NLG) correlate well with human judgement for the task of CS generation?
We propose a novel method to generate CS text from monolingual text using LLMs, which requires less data for fine-tuning. We apply this to the English-Spanish pair (RQ1). Unlike prior work, our approach leverages naturally occurring code-switched data instead of artificial examples, allowing models to learn its natural distribution. Additionally, our source texts come from informal contexts, making the output both code-switched and informal — two intertwined phenomena. We evaluate our methodology both qualitatively, with a study on human preferences and a manual error analysis, and quantitatively, using automatic NLG metrics, which allows us to study the correlation with between human and automatic evaluation for this task (RQ2).
To summarize, the contributions presented in this paper are the following:
  • we develop a novel approach for CS generation through fine-tuning LLMs on a psuedo-parallel corpus sourced from natural code-switched data.
  • we present a new parallel English-CS corpus, EN2CS, with silver train and development splits and a gold post-edited test set, which can be used to train and evaluate model for English-Spanish CS generation.
  • we perform both human and automatic evaluation of the results of models fine-tuned on this dataset, and perform a study on the correlation between human and automatic evaluation for CS generation.
@@ 2 Related Work
CS naturally occurs in communities where two or more languages are in contact, making it a subject of great interest to fields like sociolinguistics and psycholinguistics. From a social perspective, it can be affected by the attitudes of the speakers towards the languages and the CS phenomenon itself. In this respect, it is related and associated with notions of prestige and identity (Heredia et al., 2025). For example, in bilingual communities where a language is minoritized, CS can be regarded as an intrusion of the majority language (Dewaele and Wei, 2014). However, for migrant communities, it may be a way to preserve their mother tongue and as an “emblem of ethnic identity” (Poplack, 1980). Once again, its importance in different social contexts highlights the need to consider CS in NLP research, as it plays a crucial role in linguistic interactions and, consequently, the development of language technologies.
CS in NLP.
The processing and understanding of code-switched text can be crucial in the processing of social media data (Bali et al., 2014), and for speech applications, such as speech recognition or speech synthesis (Rallabandi and Black, 2017). In fact, non-monolingual speakers have shown preference for chatbots that use CS (Bawa et al., 2020). Different approaches may include normalization (Parikh and Solorio, 2021), machine translation (Xu and Yvon, 2021) or modeling code-switched text (Gonen and Goldberg, 2019). The survey by Winata et al. (2023) covers trends and advances in NLP for code-switched text, including main fields of interest and future research lines. Doğruöz et al. (2021) explain advances in applications of language technologies for code-switched text from a linguistic and social perspective.
@@
The majority of code-switched data is obtained from social media, and other popular data sources include recordings and transcriptions (Winata et al., 2023). There have been several shared tasks that deal with CS, for the tasks of Language Identification (Solorio et al., 2014; Molina et al., 2016) and Sentiment Analysis (Patwa et al., 2020). Two popular benchmarks have been created to answer the demand for evaluation of CS that covers different language pairs and tasks: LINCE (Aguilar et al., 2020), which covers traditional tasks such as Part Of Speech tagging (POS) or Sentiment Analysis (SA); and GLUECoS (Khanuja et al., 2020), which incorporates NLU tasks for the Hindi-English pair. As of today, GLUECoS cannot be used without access to the X API.
@@
CS generation has seldom been tackled in previous research. Approaches include using linguistically informed techniques that aim to find out plausible switching points (Pratapa et al., 2018; Gupta et al., 2020; Gregorius and Okadome, 2022), data augmentation (Tarunesh et al., 2021) and, more recently, prompting LLMs for CS generation (Yong et al., 2023). To the best of our knowledge, there is no previous research on CS generation with natural CS as a starting point.
@@ 3 Parallel Data Creation
In this work we present a novel approach to generate code-switched text from monolingual sentences. As a first step, we create a synthetic parallel corpus from an initial set of English-Spanish CS sentences with their English monolingual equivalents, generated by the Command R model (Cohere For AI, 2024). We exploit the fact that LLMs struggle to generate CS text given a monolingual sentence, but are able to more reliably convert a CS sentence to its corresponding monolingual version, especially when the target language is English. After having created this pseudo-parallel corpus, we use it to fine-tune LLMs on the task of conditional code-switching generation, presented in Section 4.
@@ 3.1 The LINCE benchmark
We use LINCE as a starting point, a popular benchmark that has been widely used to evaluate CS systems (Aguilar et al., 2020), which is available in 6 language pairs. This benchmark contains annotations for 5 different tasks: Language Identification (LID), Part Of Speech tagging (POS), Named Entity Recognition (NER), and Sentiment Analysis (SA). All sentences in LINCE are tokenized, and each token is annotated with a language tag as well as other categories depending on the task. In our work we focus on the English-Spanish pair and filter all sentences in the data that do not contain CS, similarly discarding all the task-specific annotations. Example 3.1 shows a random instance from LINCE.
@@
EN2CS 10,703 791 1,040
Table 1: Size of original LINCE (EN-ES) compared to the automatically filtered instances and the final set of parallel instances, dubbed EN2CS.
@@
fully enamorada de mis eyelash extensions I’m totally in love with my eyelash extensions
@@
y no te dan problemas as long as you put that it was frozen. and they don’t give you problems as long as you put that it was frozen.
Table 2: Examples of the EN2CS parallel corpus. Left: original code-switched instances, right: generated (silver) or post-edited (gold) English instances.
3.2 EN2CS
@@
We perform an initial set of experiments to determine the optimal prompt to generate monolingual English versions of the code-switched data. Ideally, we aim for a prompt that generates translations that maintain the meaning of the original sentences, are fluent and natural, whose grammar is correct and that does not contain any Spanish words or phrases. The tested prompts are listed and explained in Appendix A, and the prompt that generates outputs closest to the desired ones is: Now convert this code-switched phrase to English. Leave the parts in English as they are, focus on translating the parts in Spanish. Preliminary experiments also showed that a few-shot strategy helps the model for the task at hand, so the prompt is enriched with 5 manually selected examples that show how the output should look like.
The generated monolingual sentences are further processed, and instances that contain profanity or that are prefaced with a direct mention of the task, such as “Of course, here’s your translation:” are discarded.
In order to create a true gold standard test set to evaluate a CS generation model, we perform a manual post-edition of the the monolingual test translations for $1,040$ instances of the LINCE test set. The post-edition was carried out by three proficient speakers of English and Spanish, who were provided with specific guidelines as shown in Appendix B.
Table 1 shows the final size of the parallel corpus, which we dub EN2CS, after post-processing and post-edition, and Table 2 shows examples of silver and gold instances. The final version of our dataset therefore contains $10,703$ train and $791$ development instances with automatically translated English sentences matched to their original CS sentences, and $1,040$ gold instances with corrected English translations.
@@ 4 CS generation experiments
With EN2CS as our starting point, we frame CS generation as a machine translation task, with English as the source and CS as the target language, where parts of the source sentence have to be translated to Spanish. In our experiments we try four generative models, namely, Llama3, Llama3 Instruct Dubey et al. (2024), Mistral and Mistral Instruct Jiang et al. (2023). [^3] Llama3 models are used in their 8B size (Base and Instruct). Mistral models are used in their 7B size, version 0.3 (Base and Instruct). All the models are trained with the causal language modelling objective, but we use different input formats for the base and instruct models. For base models we follow (Zhu et al., 2024) and use templates in the form of “<X>=<Y>”, where <X> and <Y> are placeholders for the input English sentence and generated CS, respectively. At inference, the second code-switched part is left empty for the model to fill. For fine-tuning instruction-tuned models, we provide them with a system prompt that contains the instruction, a query by the user in English, and an answer from the assistant with the code-switched target. At inference time, the same system prompt is used, and the user prompt contains the English sentence, so that the model generates the assistant part. Table 3 shows examples of the format used for fine-tuning base and instruction-tuned models.
Base model
I want to not work and make money. = quiero no trabajar and make money
Instruction-tuned model
user: "I want to not work and make money."
assistant: "quiero no trabajar and make money"
Table 3: Examples and format of prompts used for finetuning base and instruction-tuned models
@@
MT Baseline damn everyone got married and we still be single lol
@@
Early experiments indicated that the models’ outputs are longer than expected and usually produce the desired output up to a punctuation mark and then either begin to translate the sentence again or hallucinate more content. We therefore truncate the output up to a punctuation mark where the length is closest to that of the original sentence. We additionally experimented with the length_penalty and exponential_length_decay generation parameters, as well as trying to control the length of the generation with length codes, but find that the truncation heuristic performs the best. Accordingly, all further experiments will use the truncated output. This overgeneration problem has been reported in previous papers, where similar truncation strategies have been adopted (Bawden and Yvon, 2023).
We also include a dedicated encoder-decoder model as a baseline, trained on EN2CS using the MarianNMT toolkit (Junczys-Dowmunt et al., 2018), with a mini-batch size automatically selected for the available memory [^4] We used two GPUs and $8GB$ per GPU.. Optimization is performed using Adam (Kingma and Ba, 2015), with $\alpha=0.0003$, $\beta_{1}=0.9$, $\beta_{2}=0.98$ and $\epsilon=10^{-9}$, using a standard learning rate of $3e^{-4}$. Validation is conducted every $100$ steps, and training stops if perplexity showed no improvement after 5 consecutive checkpoints.
Table 4 shows an example of the outputs of the different models, compared to the original code-switched sentence and the English monolingual sentence that they received as input.
@@ 5.1 Preference based evaluation
We perform a tournament-based evaluation that allows us to determine the ranking of models in terms of human preference. A total of $660$ instances are matched against each other, corresponding to the outputs of the five models for $110$ English source sentences, as well as the gold standard reference. The evaluation is conducted pairwise, requiring annotators to choose the best out of two sentences or declare a tie. When choosing the best sentence, annotators do not know the original English sentence, nor which model produced what output. This process results in 110($650110\cdot{6\choose 2}=1,$ 650 comparisons, and was carried out by $11$ annotators, with each annotator performing $150$ random comparisons.
Annotators are provided with a series of criteria to choose between the instances, devised after the error analysis described in the next section. They must take into account three main criteria, that must be applied in the following order: a) the presence and naturalness of the CS; b) the content and fluency of the sentences; and c) the orthographical errors of the instances (correct punctuation, presence of typos, etc.). Annotators are furthermore asked to avoid declaring ties, unless completely necessary (e.g., in a case where both sentences are completely monolingual and therefore equally incorrect), to compel them to develop a preference. The complete annotation guidelines are available in Appendix C.
Ranking Model Score
[Uncaptioned image] Gold Standard 392.5
[Uncaptioned image] Llama3 325.5
[Uncaptioned image] Llama3 Instruct 303.0
4 Mistral 285.5
5 Mistral Instruct 242.0
6 MT Baseline 101.5
Table 5: Ranking of models according to the human preference score.
With the results of the pairwise comparisons, we calculate a global score for each model, as follows: every time a model is voted, it gets $1$ point, and the loser gets 00 points; in case of ties, both models get $0.5$ points each. Table 5 shows the global scores, as well as the ranking of human preferences according to said score. We find that the gold standard reference obtains the highest score, as expected, and that Llama3 ranks the highest among the automatic methods. Instruction-tuned models obtain worse scores compared to their base model counterparts, with a similar difference for both Llama3 and Mistral family of models. Finally, the evaluation also shows that humans clearly prefer the output of systems based on Large Language Models compared to the MT baseline, which is voted as the worst model by a large margin.

### v2

@@ Abstract
Code-switching (CS) is still a critical challenge in Natural Language Processing (NLP), due to the limited availability of large-scale, diverse CS datasets for robust training and evaluation. Despite recent advances, the true capabilities and limitations of LLMs in handling CS remain underexplored. In this work, we investigate the extent to which LLMs can be used in a framework for CS text generation, focusing on the English-Spanish language pair. Our proposed methodology consists of back-translating natural CS sentences into monolingual English, and using the resulting parallel corpus to fine-tune LLMs to turn monolingual sentences into CS. We thoroughly analyse the models’ performance through a study on human preferences, a qualitative error analysis, an evaluation with popular reference-based metrics and LLM-based judgment. Results show that fine-tuning can be a key step to ensure that current LLMs consistently generate fluent code-switched text and that our methodology generates high-quality outputs, expanding research opportunities in CS communication. We find that traditional metrics do not correlate with human judgement when assessing the quality of the generated CS data but LLM-based judgment aligns more closely with human preferences. We release our code and generated dataset under a CC-BY-NC-SA license. [^1] URL to be announced upon acceptance.
@@ 1 Introduction
Despite the prevalence of code-switching, most research in Natural Language Processing (NLP) assumes monolingualism as a standard for human communication. However, this implicit decision means that state-of-the-art models are not able to properly interpret or generate CS data. Even advances in multilingual language modelling (Lin et al., 2022; Chowdhery et al., 2023) have not led to significant improvements, and performance on CS data is still poor compared to performance on monolingual data (Aguilar et al., 2020; Winata et al., 2021). This occurs because there is little CS text available in the multilingual pretraining data. Similarly, there are no parallel datasets available to learn to generate CS in a supervised fashion, as one would expect for tasks such as machine translation. Finally, existing methodology for evaluating automatically generated CS text, which has specific needs different from other text generation tasks, are still not good enough and fail to capture nuances of CS text (Srivastava and Singh, 2021). It is therefore crucial to develop methodologies to enable models to generate natural CS text and simultaneously implement robust evaluation frameworks that can assess how well NLP systems handle CS across multiple tasks. We argue that both of these goals require models that can conditionally generate CS from monolingual text. Consequently, our research focuses on the development of a methodology to fine-tune and evaluate LLMs on the task of CS generation, following three main research questions:
@@
What are the comparative strengths and limitations of fine-tuned versus non-fine-tuned LLMs in generating fluent and natural code-switched text?
@@
How can we leverage LLMs to create high-quality pseudo-parallel data for fine-tuning LLMs in CS text generation?
RQ3:
Do automatic metrics for Natural Language Generation (NLG) or LLM judges correlate well with human judgement for the task of CS generation?
Based on these research questions, we propose a novel approach to generate CS from monolingual text using LLMs and apply it to the English-Spanish pair. We create a new parallel English–CS corpus, EN-CS, by leveraging natural CS data and using LLMs to perform back-translation from CS into English, resulting in high-quality pseudo-parallel pairs, suitable for training and evaluating models on CS generation (RQ2). We provide a comprehensive comparison of CS generation using LLMs in both zero-shot and fine-tuned settings, and we compare their performance against that of a dedicated machine translation (MT) model (RQ1). Finally, we evaluate our methodology both qualitatively, with a study on human preferences and a manual error analysis, and quantitatively, using automatic NLG metrics and LLM as a judge, which allows us to study the correlation between human and automatic evaluation for this task (RQ3).
@@ 2 Related Work
CS naturally occurs in communities where two or more languages are in contact, making it a subject of interest to fields like sociolinguistics and psycholinguistics. From a social perspective, it can be affected by speakers’ attitudes towards the languages and the CS phenomenon itself. In this respect, it is related to notions of prestige and identity (Heredia et al., 2025). For example, in bilingual communities where a language is minoritized, CS can be seen as an intrusion of the majority language (Dewaele and Wei, 2014). However, for migrant communities, it may be a way to preserve their mother tongue and as an “emblem of ethnic identity” (Poplack, 1980). Its importance in different social contexts highlights the need to consider CS in NLP research, as it plays a crucial role in linguistic interactions and, consequently, the development of language technologies.
@@
Most code-switched data stems from social media, while other popular data sources include recordings and transcriptions (Winata et al., 2023). Shared tasks using such CS data have been organized for the tasks of Language Identification (Solorio et al., 2014; Molina et al., 2016) and Sentiment Analysis (Patwa et al., 2020). Similarly, two benchmarks exist to evaluate model performance on CS text, covering different language pairs and tasks: LINCE (Aguilar et al., 2020), which covers tasks such as Part Of Speech tagging or Sentiment Analysis; and GLUECoS (Khanuja et al., 2020), which focuses on NLU tasks for Hindi-English. Unfortunately, GLUECoS cannot be currently used without access to the X API.
@@
CS generation has seldom been tackled in previous research. Approaches include linguistically informed techniques to find plausible switching points (Pratapa et al., 2018; Gupta et al., 2020; Gregorius and Okadome, 2022; Hsu et al., 2023), data augmentation (Tarunesh et al., 2021) and, more recently, prompting LLMs for CS generation (Yong et al., 2023). While CS generation is often evaluated by human annotators (Tarunesh et al., 2021; Gregorius and Okadome, 2022), there remains a need for robust automatic evaluation methodologies to assess the naturalness and fluency of the generated texts, with some recent studies already exploring approaches like Judge-LLMs (Kuwanto et al., 2024).
@@ 3 Parallel Data Creation
In this work we present a novel approach to generate code-switched text from monolingual sentences. As a first step, we create a synthetic parallel corpus from an initial set of English-Spanish CS sentences from the LINCE benchmark (Aguilar et al., 2020) with their English monolingual equivalents, generated by the Command R model (Cohere For AI, 2024). We exploit the fact that LLMs struggle to generate CS text given a monolingual sentence (c.f. Section 5), but are able to more reliably convert a CS sentence to its corresponding monolingual version, especially when the target language is English. After having created this pseudo-parallel corpus, we use it to fine-tune LLMs on the task of conditional code-switching generation, presented in Section 4.
@@ 3.1 The LINCE benchmark
We use LINCE as a starting point, a popular benchmark that has been widely used to evaluate CS systems (Aguilar et al., 2020), which is available in 6 language pairs. All sentences in LINCE are tokenized, and each token is annotated with a language tag as well as other categories depending on the task. In our work we focus on the English-Spanish pair and filter all sentences in the data that do not contain CS, similarly discarding all the task-specific annotations. Example 3.1 shows a random instance from LINCE.
@@
EN-CS 10,703 791 1,040
Table 1: Size of original LINCE (EN-ES) compared to the automatically filtered instances and the final set of parallel instances, dubbed EN-CS.
@@
Table 2: Examples of the EN-CS parallel corpus. Left: original code-switched instances, right: generated (silver) or post-edited (gold) English instances.
3.2 EN-CS
@@
We perform an initial set of experiments to determine the optimal prompt to generate monolingual English versions of the code-switched data. Ideally, we aim for a prompt that generates translations that maintain the meaning of the original sentences, are fluent and natural, whose grammar is correct, and that do not contain any Spanish words or phrases. After extensive testing (see Appendix A), we use the following prompt in a 5-shot setting: Now convert this code-switched phrase to English. Leave the parts in English as they are, focus on translating the parts in Spanish. Finally, we filter output instances that contain profanity that was not present in the source texts or irrelevant information, such as “Of course, here’s your translation:”.
In order to create a valid gold standard test set, we perform a manual post-edition of the the monolingual test translations for $1,040$ instances of the LINCE test set. The post-edition was carried out by three proficient speakers of English and Spanish, who were provided with specific guidelines as shown in Appendix B.
Table 1 shows the final size of the parallel corpus, which we dub EN-CS, after post-processing and post-edition, and Table 2 shows examples of silver and gold instances. The final version of our dataset therefore contains $10,703$ train and $791$ development instances with automatically translated English sentences matched to their original CS sentences, and $1,040$ gold instances with post-edited English translations.
3.2.1 Quality assessment
We evaluate the quality of the automatic translations (train/dev) by measuring two dimensions: the overall fluency of the sentences and adequacy of the translations in respect to the source texts. Two fluent English-speaking annotators evaluate the same 100 random instances using a 5-point Likert scale Callison-Burch et al. (2007) and obtain 4.6 (fluency) and 4.5 (adequacy) points on average, which show the quality of the generated translations. A quadratic Cohen’s $\kappa$ of $0.57$ indicates moderate agreement, likely due to lower Likert scores (1, 2, and 3) being rarely selected by the annotators, which is a known problem for $\kappa$ Xu and Lorber (2014); Barnes et al. (2025). In fact, the raw agreement between annotators is substantially higher: 0.71 for fluency and 0.65 for adequacy. See Appendix C for further details.
To estimate the quality of the post-edition process, we compare the post-editions of two annotators on 100 additional random instances. The results show that 75% of the sentences remain unchanged, as they are already adequate. There is a 87.87% similarity between the post-editions of the two annotators, as measured by Levenshtein distance, demonstrating a high degree of consistency and quality in the post-edition process.
@@ 4 CS generation experiments
With EN-CS as our starting point, we frame CS generation as a machine translation task, with English as the source and CS as the target language, where parts of the source sentence have to be translated to Spanish. In our experiments, we fine-tune four small-sized generative models, namely, Llama3 8B, Llama3 Instruct 8B Dubey et al. (2024), Mistral 7B and Mistral Instruct 7B Jiang et al. (2023).
To fine-tune the models, we use the causal language modelling objective, but with appropriate input formats for the base and instruct models. For base models we use templates (Zhu et al., 2024) in the form of “<X>=<Y>”, where <X> and <Y> are placeholders for the input English sentence and generated CS, respectively. At inference, the second code-switched part is left empty for the model to fill. For instruction-tuned models, we provide a system prompt with the instruction, a query by the user in English, and an answer from the assistant with the code-switched target. At inference time, the answer is left blank (See Table 7 in Appendix D for example prompts).
@@
Llama3.3-70Bfs damn todo el mundo se casó y nosotros seguimos solteros lol forever alone
GPT-4ofs damn todos se casaron y nosotros seguimos solteros lol forever alone
NLLB damn everyone got marry and its still single lol forever alone
@@
Early experiments indicated that fine-tuned models usually produce the desired output up to a punctuation mark and then either begin to translate the sentence again or hallucinate more content. We therefore truncate the output up to a punctuation mark where the length is closest to that of the original sentence (Bawden and Yvon, 2023).
As baselines, we include few-shot experiments by directly prompting GPT-4o and Llama3.3 70B Instruct to generate CS text using a 5-shot approach (see Table 7 in the Appendix). We refer to these systems in the experiments as GPT-4ofs and Llama3.3-70Bfs, respectively. We also include a strong dedicated MT baseline, developed by fine-tuning the NLLB team et al. (2022) model (nllb-200-distilled-600M) using EN-CS. The model was trained with standard settings.  [^3] A batch size of 32, learning rate of $1\times 10^{-4}$, using constant learning rate schedule with 1,000 warmup steps, a gradient clipping threshold of 1.0, and a weight decay of $1\times 10^{-3}$. Training was conducted for 50,000 steps. For evaluation, we selected the checkpoint that achieved the highest BLEU score on the development set.
Table 3 shows an example of the outputs of the different models, compared to the original code-switched sentence, and the English monolingual sentence that they received as input.
@@ 5.1 Preference based evaluation
We perform a tournament-based evaluation that allows us to determine the ranking of models in terms of human preference. A total of $880$ instances are matched against each other, corresponding to the outputs of the seven models for $110$ English source sentences, as well as the gold standard reference. The evaluation is conducted pairwise, requiring annotators to choose the best out of two sentences or declare a tie. When choosing the best sentence, annotators do not know the original English sentence, nor which model produced what output. This process results in $080110\cdot{8\choose 2}=3,$ 080 comparisons, and was carried out by $14$ annotators, with each annotator performing at most $300$ random comparisons.
Annotators are provided with a series of criteria to choose between the instances, based on the error analysis described in the next section. They must take into account three main criteria, which must be applied in the following order: a) the presence and naturalness of the CS; b) the content and fluency of the sentences; and c) the orthographical errors of the instances (correct punctuation, presence of typos, etc.). Annotators are furthermore asked to avoid declaring ties, unless completely necessary (e.g., in a case where both sentences are completely monolingual and therefore equally incorrect) to compel them to develop a preference. The complete annotation guidelines are available in Appendix E. Inter-annotator agreement on a subset of 100 sentence pairs shows substantial agreement ($\kappa=0.74$).
Human GPT
Model Score Rank Score Rank
Gold Standard 525.0 [Uncaptioned image] 484.5 [Uncaptioned image]
Llama3 423.0 [Uncaptioned image] 374.0 4
Llama3 Instruct 393.0 [Uncaptioned image] 367.5 5
NLLB 391.0 4 304.0 8
Mistral 368.0 5 334.5 6
GPT-4ofs 366.5 6 478.0 [Uncaptioned image]
Llama3.3-70Bfs 307.0 7 428.0 [Uncaptioned image]
Mistral Instruct 306.5 8 309.5 7
Table 4: Ranking of models according to human preference (c.f. Section 5.1) and using GPT as a judge (c.f. Section 6.2).
We calculate a global score for each model, as follows: every time a model is voted, it gets $1$ point, and the loser gets 00 points; in case of ties, both models get $0.5$ points each. Table 4 shows the global scores, as well as the ranking of human preferences according to said score (second and third columns). We find that the gold standard reference obtains the highest score, as expected, and that fine-tuned Llama3 ranks the highest among the automatic methods. Instruction-tuned models obtain worse scores compared to their base model counterparts, with a similar difference for both Llama3 and Mistral family of models. The NLLB model ranks higher than the Mistral models, but lower than the fine-tuned Llama3 models, showing their potential to rival dedicated models in generation tasks. According to these preferences, fine-tuning LLMs for CS generation can be critical to ensure better results, since the larger models with few-shot prompting rank only above Mistral Instruct, with GPT-4ofs outranking Llama3.3-10Bfs.